OpenStudy (anonymous):

Cora is graphing the equation y = 12x. She has placed a point on the origin and uses the slope to find 2 more points. Which points does Cora plot on her graph? Choose exactly 2 answers that are correct. A. (-1, -12) B. (-12, -1) C. (1, 12) D. (12, 1)

OpenStudy (compassionate):

So, y = 12x... This is simpler than a lot of people make it. Check this out. We know points are in the form (x, y), and we have y = mx + b So... y = 12x -- right? Lets take the first one, A: (-1, -12) Remember.. (x, y) (-1, -12) x = -1 y = -12 -------- y = 12x -12 = 12(-1) -12 = -12... WHICH IS TRUE! So one of your points is A. Find the other doing this method. :)

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (compassionate):

Do you understand, Nichole? It's really simple! Just substitute (x, y) into your y = mx + b equation. Say like; you get 10 = 10 This is TRUE and solution But if you get 3 = 5 This is FALSE and you need to try a new pair of (x, y) Take care! :)

OpenStudy (anonymous):

Yes I understand thank for explaining :)

OpenStudy (anonymous):

Thanx*